# Dark Matter (DM); Dark Stars (DSs); High-z galaxies; Supermassive Black Holes (SMBHs), Little Red Dots (LRDs); Blue Monsters; Over-massive BH galaxies (OBGs)

### v1

@@ 5.1 Mergers
UHZ1’s morphology is “evocative of late stage mergers at low redshifts” (Bogdan:2023UHZ1). As such, the merger of two DM halos is the most plausible explanation of UHZ1. We discuss this scenario below, in the context of a Dark Star (DS) seed for the SMBH powering the X-ray spectra of UHZ1. A DS seed forms in a DM halo at $z_{form}\in[15-25]$ (see Sec. 4). Subsequently, the halo hosting the (by now) SMDS begins to merge with another halo (either dark or hosting nuclear fusion-powered stars). The SMDS, due to its large mass, may remain at the center of the new larger object from the merger and continue to grow. However, it is also possible that, as a result of the merger, sometimes the SMDS gets slightly dislodged from the high DM density at the center of the DM halos, and as such experience a sudden and destabilizing decrease in the DM heating.  [^7] For details see Appendix D.
In addition to the collapse of the SMDS to a SMBH, the merger could replenish the baryonic reservoir and trigger a starburst episode in the vicinity of the SMBH (e.g. Silk:2005; Levin:2007MNRAS; carilli:2013; heckman:2014; MF2023). Thus, the SMBH can continue to grow at the Eddington rate, and be observed as the X-ray portion of the UHZ1 radiation. The stellar population embedded in those two merging halos is responsible for the majority of the IR observed data. Alternatively, UHZ1 can be explained by a halo hosting a SMBH seeded by a DS merges with another halo, leading to a starburst episode in the newly formed system. In Fig. 7 we present a schematic representation of our DS model for UHz1 discussed here.
@@ 5.2 Same DM Halo Host
Photons in the the Lyman-Werner (LW) band (11.213.611.2-13.6 eV) can photo-dissociate molecular H2 inside star forming H clouds, leading to a quenching of the star formation rate. Omukai:1999 shows that a Pop III star hotter than $T_{eff}\gtrsim 54,000$ K (i.e. with mass $M\gtrsim 40M_{\odot}$) can inhibit star formation in a region surrounding it as large as $r_{sh}\sim 1$kpc. Compared to the size of a high redshift ($z\sim 20$) DM halo, 1 kpc can be significant, or sometimes even exceed it. This would imply that star formation is almost turned off in any high redshift DM halo, as soon as the first massive Pop III star(s) are born within it and if there are no other efficient cooling mechanisms besides H2.
@@
Ref. Omukai:1999 makes two key simplifying assumptions: i) equilibrium between the dissociation and formation rates of H2 within the vicinity of the star is quickly reached, and ii) the baryonic number density is constant throughout the DM halo. Accounting for the time required for the LW flux of a massive Pop III star to photo-dissociate H2 within its surroundings allows for star formation to continue efficiently in the most massive H2 cooling DM halos ($M\sim 10^{7}M_{\odot}$) or in the densest baryonic clumps (independent of the host halo mass) (Glover:2000). Moreover, if assumption (ii) is replaced with a commonly used isothermal sphere profile for the gas density surrounding the star (or any other realistic profile) the estimates of the radius of influence $r_{sh}$ would be significantly reduced.
@@
Dark Stars are inherently cooler than Pop III stars, so their radius of influence would be significantly smaller. A SMDS of $\sim 10^{4}M_{\odot}$ emits the same LW flux as the $40M_{\odot}$ Pop III star considered by Omukai:1999; Glover:2000. Therefore it can take up to $10$ Myrs (assuming an accretion rate of $10^{-3}M_{\odot}\mathrm{yr}^{-1}$) from the moment a DS forms until it begins to significantly suppress the star formation in its host halo. As such, it is entirely possible that Pop III stars can form from the collapse of gas $H_{2}$ clumps located the same parent DM halo as a Dark Star growing towards a SMDS. Once any of those Pop III stars grows sufficiently massive the LW radiation will inhibit other stars from forming in their parent DM halo. Without mergers, in this scenario the total stellar mass is not expected to exceed $\sim 10^{6}M_{\odot}$, in view of the fact that a higher mass would imply $M_{halo}\gtrsim 10^{7}M_{\odot}$, i.e. an atomically cooled halo. We discuss this scenario next.
@@
For DM halos with a virial temperature $T_{vir}\gtrsim 10^{4}$ K, atomic hydrogen cooling becomes efficient. As such, gas can cool irrespective of the flux of the LW radiation and regular star formation can continue indefinitely. Atomic Cooling Halos (ACHs) have a mass greater than $\sim 10^{7}M_{\odot}$. The formation of zero metallicity clusters at the center of ACHs has been studied in simulations, in the context of Population III stars (e.g. Kimm:2016clusers; liu:2024universal). The growth of SMDSs at the centers of atomic cooling halos was studied using the MESA stellar evolution code in Rindler-Daller:2014uja. Since the baryonic mass within the UHZ1 system must be greater than $10^{7}M_{\odot}$, in the absence of halo mergers discussed in Sec. 5.1, the host DM halo must have been an atomic cooling halo. Thus it is plausible (although unlikely, in view of UHZ1’s distorted morphology) that all the stars within UHZ1 (including the Supermassive Dark Star we assume as the seed for the SMBH) have formed within the same DM halo.
@@ 6 Discussion and Conclusions
One of the challenges of the “low mass seed" scenarios (in which a quasar at $z\gtrsim 6$ originates from sustained Eddington-rate accretion onto an initially few stellar–mass BH seeded by a Pop III star) is feedback. This can easily unbound the gas, thus arresting the growth of the BH. Furthermore, growing small BHs through mergers is also challenging since the rate is just too slow. Moreover, gravitational recoils during BH mergers can cause them to be ejected from the original halos (Whalen:2004; Haiman:2001; Haiman:2004; Whalen:2020). Finally, recent X-ray to NIR observations of high redshift QSOs suggest that the intrinsic photon index from 0.5 to 10 keV cannot be explained by high Eddington ratios alone (farina:2022). These lines of evidence all favor a “heavy seed" scenario, as further demonstrated by Bogdan:2023UHZ1. The IR luminosities, obscurations, and general observations of massive galaxies at the highest redshift suggest that the heavy seed black holes must grow without severely impacting star formation and dust production in the halos. Fueled by Dark Matter annihilations favored by the physical conditions in high redshift ($z\sim[10-25]$) Dark Matter halos, Dark Stars provide a solution to both enigmas of large galaxies and luminous QSOs observed by JWST at the highest redshifts.
@@
The abundance of halos at $z\gtrsim 10$ predicted to host DCBHs is small, according to simulations (e.g. DCBHsDensity). Thus, once a larger sample of $z\gtrsim 10$ SMBHs is observed, one could place constraints on the DCBH scenario. While DSs have not yet been realized in cosmo-hydrodynamical simulations, we expect their density, at $z\gtrsim 10$ to be greater than that of DCBHs, because
@@
    DSs can form in both ACHs and $H_{2}$ cooling halos (Spolyar:2008dark).
@@
    Since gas fragmentation is directly suppressed by DM annihlations (Stacy:2014), DSs formation does not require a mechanism to destroy $H_{2}$.
@@
    Because they are naturally puffy (Spolyar:2008dark; Spolyar:2009; Freese:2010smds), DSs can form even when baryonic infall is low. In contrast, DCBHs can form only if the accretion rate onto the Pop III progenitor is $\gtrsim 0.1M_{\odot}$ (Hosokawa_2012; Toyouchi_2022).
@@
As discussed in Sec. 3, SMDSs can provide a solution to the enigmatic Blue Monsters uncovered by JWST. Those are extremely compact ($r_{e}\lesssim 400$p) yet very bright objects at $z\gtrsim 10$, which show very little dust attenuation and, if interpreted as galaxies have to have converted gas to stars at incredibly high rates. Compactness, lack of dust attenuation, and high conversion rate of gas to stellar material are, in fact, a feature of the SMDS model, thus turning an astronomical puzzle into a potential line of evidence for SMDSs. Furthermore, We argue that Dark Stars can also be part of the solution to the Little Red Dots (LDRs) puzzle (e.g. LDRs:2024ApJ...963..129M; LRDsCensus:2024ApJ...968...38K; LRDs:2025ApJ...991...37A). Those are extremely compact sources unlikely to be galaxies due to incredibly dense stellar packing and little to no obscuration due to dust (LDRsNoALMA:2025ApJ...990L..61C). Instead many LRDs seem to point towards the existence of SMBHs surrounded by remnant stellar layers that explain the weak UV and lack of X-ray emissions (e.g. LRDsBHStar:2025arXiv250316596N; LRDBHStar:2025A&A...701A.168D; LRDsBHStars:2025arXiv250709085B). Such objects, a.k.a. Black Hole Stars or “quasi-stars” are hypothesized in the DCBH scenario (Belgman:2006) and in future work we plan to investigate if such quasi-stars they can also form via the collapse of SMDSs. In addition, a large sub-sample of JWST’s Little Red Dots (LRDs) are thought to harbor BHs that are significantly overmassive relative to their host galaxies (LRDsOvermassiveBHs:2025ApJ...985..169D). Each of those Over-Massive Black Hole Galaxies (OBGs) can be easily explained by Dark Stars via the one of the two mechanisms we proposed in this paper for UHZ1: mergers (Sec. 5.1) or same halo (Sec. 5.2).
Moreover, Dark Stars could affect the re-ionization history of the Universe (Scott_2011). As shown by Gondolo_2022, the inclusion of a population of DSs can naturally lead values of CMB electron scattering optical depth as low as $\tau\sim 0.05$, in agreement with the Planck measurement. This value is impossible to reconcile with only a Pop III/II stellar population responsible for reionization, unless either very low star formation efficiencies, or strong LW feedback are invoked. In contrast, in view of their small predicted abundance, Direct Collapse Black Holes are not expected to contribute significantly to re-ionization history or to its CMB markers.
@@ Appendix A Dark Matter, a brief overview
Since Dark Stars rely on Dark Matter particles as a power engine, or, conversely, as they can be used as DM probes, we briefly review here the main evidence for the existence of Dark Matter (DM) and strategies for its eventual discovery. The fundamental nature of DM has remained largely elusive since Fritz Zwicky coined the term Dunkle Materie (i.e. Dark Matter) in 1933 (Zwicky:1933). Its gravitational effects have since been inferred experimentally at scales ranging from galactic (e.g. Rubin:1970), galactic cluster (e.g. Natarajan:2017), and cosmological (e.g. Madhavacheril:2014; Vikram:2015; Hikage:2018; Aghanim:2018). Moreover, DM plays a significant role in the early universe, leaving its imprint in afterglow radiation from the Big Bang, the so-called Cosmic Microwave Background (CMB) radiation. A combination of all those data leads to the emergence of the concordance cosmological model ($\Lambda$-CDM), where the vast majority of the energy budget of the Universe is in the form of two mysterious components: Dark Energy ($70\%$) and Dark Matter (25%). The former is responsible for the current accelerated expansion of our universe (Riess:1998SN), while the latter assembles in large scale filamental structures, thus providing a scaffolding for the ordinary matter to assemble in galaxies and clusters. The thought that all the stars in all the galaxies in our universe form less than one percent of the total energy budget of our Universe is unsettling, perhaps even scary.
@@
It is now firmly established that the majority of the DM in the Universe comes in the form of new, yet to be discovered particles. Amongst those, Axions and Weakly Interacting Massive Particles (WIMPs) are perhaps the two leading candidates, in view of their natural theoretical motivations. There are three different strategies with which one can approach the hunt for Dark Matter particles: direct detection experiments (XENON (e.g. Aprile:2018; Aprile:2020), LZ (e.g. LZ:2022lsv; LZ:2024psa), PICO (e.g. Amole:2019fdf), etc), particle production in accelerators (e.g. boveia2018dark) and lastly, indirect detection. The later method relies on potential observable signals of DM self-annihilating, or decaying, in various astrophysical objects or environments. As of yet, none of those methods lead to a discovery signal, although there are tantalizing hints, especially from indirect detection methods. For example, both the AMS-02 anomalous anti-proton signal, and the Fermi Gamma ray excess from the Galactic Center, can both be explained by the self annihilation of $\sim 60$ GeV WIMPs (Cholis:2019). For both of those excesses there are also non-DM astrophysical interpretations. This is possible because of the somewhat large uncertainties, at a theoretical level, in the backgrounds for both of those signals. For reviews on Dark Matter and its theoretical or observational status the reader can consult Refs. (Freese:2017dm; arbey2021dark).
@@
We end this section with a non exhaustive list of objects that can act as indirect DM probes: neutron stars (e.g. Bramante:2017; Baryakhtar:2017; Bell:2018pkk; Bell:2020; Ilie:2020Comment; Nguyen:2022NS), white dwarfs (e.g. Dasgupta:2019juq; Horowitz:2020axx), Jovian planets (e.g. Leane:2020wob; Croon:2023DMExoForm), present day stars capturing and burning Dark Matter (e.g. Moskalenko:2007ApJ; Scott:2009DS; John:2024thz), and, of particular interest for this work, the first generation of stars in the Universe (e.g. Freese:2008cap; Iocco:2008cap; Ilie:2020PopIII; Ilie:2021mcms).

### v2

@@ 5.1 Mergers
UHZ1’s morphology is “evocative of late stage mergers at low redshifts” (Bogdán et al., 2024). As such, the merger of two DM halos is the most plausible explanation of UHZ1. We discuss this scenario below, in the context of a Dark Star (DS) seed for the SMBH powering the X-ray spectra of UHZ1. A DS seed forms in a DM halo at $z_{form}\in[15-25]$ (see Sec. 4). Subsequently, the halo hosting the (by now) SMDS begins to merge with another halo (either dark or hosting nuclear fusion-powered stars). The SMDS, due to its large mass, may remain at the center of the new larger object from the merger and continue to grow. However, it is also possible that, as a result of the merger, sometimes the SMDS gets slightly dislodged from the high DM density at the center of the DM halos, and as such experience a sudden and destabilizing decrease in the DM heating.  [^8] For details see Appendix D.
In addition to the collapse of the SMDS to a SMBH, the merger could replenish the baryonic reservoir and trigger a starburst episode in the vicinity of the SMBH (e.g. Silk, 2005; Levin, 2007; Carilli and Walter, 2013; Heckman and Best, 2014; Mercedes-Feliz et al., 2023). Thus, the SMBH can continue to grow at the Eddington rate, and be observed as the X-ray portion of the UHZ1 radiation. The stellar population embedded in those two merging halos is responsible for the majority of the IR observed data. Alternatively, UHZ1 can be explained by a halo hosting a SMBH seeded by a DS merges with another halo, leading to a starburst episode in the newly formed system. In Fig. 7 we present a schematic representation of our DS model for UHz1 discussed here.
@@ 5.2 Same DM Halo Host
Photons in the the Lyman-Werner (LW) band (11.213.611.2-13.6 eV) can photo-dissociate molecular H2 inside star forming H clouds, leading to a quenching of the star formation rate. Omukai and Nishi (1999) shows that a Pop III star hotter than $T_{eff}\gtrsim 54,000$ K (i.e. with mass $M\gtrsim 40M_{\odot}$) can inhibit star formation in a region surrounding it as large as $r_{sh}\sim 1$kpc. Compared to the size of a high redshift ($z\sim 20$) DM halo, 1 kpc can be significant, or sometimes even exceed it. This would imply that star formation is almost turned off in any high redshift DM halo, as soon as the first massive Pop III star(s) are born within it and if there are no other efficient cooling mechanisms besides H2.
@@
Ref. Omukai and Nishi (1999) makes two key simplifying assumptions: i) equilibrium between the dissociation and formation rates of H2 within the vicinity of the star is quickly reached, and ii) the baryonic number density is constant throughout the DM halo. Accounting for the time required for the LW flux of a massive Pop III star to photo-dissociate H2 within its surroundings allows for star formation to continue efficiently in the most massive H2 cooling DM halos ($M\sim 10^{7}M_{\odot}$) or in the densest baryonic clumps (independent of the host halo mass) (Glover and Brand, 2001). Moreover, if assumption (ii) is replaced with a commonly used isothermal sphere profile for the gas density surrounding the star (or any other realistic profile) the estimates of the radius of influence $r_{sh}$ would be significantly reduced.
@@
Dark Stars are inherently cooler than Pop III stars, so their radius of influence would be significantly smaller. A SMDS of $\sim 10^{4}M_{\odot}$ emits the same LW flux as the $40M_{\odot}$ Pop III star considered by Omukai and Nishi (1999); Glover and Brand (2001). Therefore it can take up to $10$ Myrs (assuming an accretion rate of $10^{-3}M_{\odot}\mathrm{yr}^{-1}$) from the moment a DS forms until it begins to significantly suppress the star formation in its host halo. As such, it is entirely possible that Pop III stars can form from the collapse of gas $H_{2}$ clumps located the same parent DM halo as a Dark Star growing towards a SMDS. Once any of those Pop III stars grows sufficiently massive the LW radiation will inhibit other stars from forming in their parent DM halo. Without mergers, in this scenario the total stellar mass is not expected to exceed $\sim 10^{6}M_{\odot}$, in view of the fact that a higher mass would imply $M_{halo}\gtrsim 10^{7}M_{\odot}$, i.e. an atomically cooled halo. We discuss this scenario next.
@@
For DM halos with a virial temperature $T_{vir}\gtrsim 10^{4}$ K, atomic hydrogen cooling becomes efficient. As such, gas can cool irrespective of the flux of the LW radiation and regular star formation can continue indefinitely. Atomic Cooling Halos (ACHs) have a mass greater than $\sim 10^{7}M_{\odot}$. The formation of zero metallicity clusters at the center of ACHs has been studied in simulations, in the context of Population III stars (e.g. Kimm et al., 2016; Liu et al., 2024). The growth of SMDSs at the centers of atomic cooling halos was studied using the MESA stellar evolution code in Rindler-Daller et al. (2015). Since the baryonic mass within the UHZ1 system must be greater than $10^{7}M_{\odot}$, in the absence of halo mergers discussed in Sec. 5.1, the host DM halo must have been an atomic cooling halo. Thus it is plausible (although unlikely, in view of UHZ1’s distorted morphology) that all the stars within UHZ1 (including the Supermassive Dark Star we assume as the seed for the SMBH) have formed within the same DM halo.
@@ 6 Discussion and Conclusions
One of the challenges of the “low mass seed" scenarios (in which a quasar at $z\gtrsim 6$ originates from sustained Eddington-rate accretion onto an initially few stellar–mass BH seeded by a Pop III star) is feedback. This can easily unbound the gas, thus arresting the growth of the BH. Furthermore, growing small BHs through mergers is also challenging since the rate is just too slow. Moreover, gravitational recoils during BH mergers can cause them to be ejected from the original halos (Whalen et al., 2004; Haiman and Loeb, 2001; Haiman, 2004; Whalen et al., 2020). Finally, recent X-ray to NIR observations of high redshift QSOs suggest that the intrinsic photon index from 0.5 to 10 keV cannot be explained by high Eddington ratios alone (Farina et al., 2022). These lines of evidence all favor a “heavy seed" scenario, as further demonstrated by Bogdán et al. (2024). The IR luminosities, obscurations, and general observations of massive galaxies at the highest redshift suggest that the heavy seed black holes must grow without severely impacting star formation and dust production in the halos. Fueled by Dark Matter annihilations favored by the physical conditions in high redshift ($z\sim[10-25]$) Dark Matter halos, Dark Stars provide a solution to both enigmas of large galaxies and luminous QSOs observed by JWST at the highest redshifts.
@@
The abundance of halos at $z\gtrsim 10$ predicted to host DCBHs is small, according to simulations (e.g. Habouzit et al., 2016). Thus, once a larger sample of $z\gtrsim 10$ SMBHs is observed, one could place constraints on the DCBH scenario. While DSs have not yet been realized in cosmo-hydrodynamical simulations, we expect their density, at $z\gtrsim 10$ to be greater than that of DCBHs, because
@@
    DSs can form in both ACHs and $H_{2}$ cooling halos (Spolyar et al., 2008).
@@
    Since gas fragmentation is directly suppressed by DM annihlations (Stacy et al., 2014), DSs formation does not require a mechanism to destroy $H_{2}$.
@@
    Because they are naturally puffy (Spolyar et al., 2008, 2009; Freese et al., 2010), DSs can form even when baryonic infall is low. In contrast, DCBHs can form only if the accretion rate onto the Pop III progenitor is $\gtrsim 0.1M_{\odot}$ (Hosokawa et al., 2012; Toyouchi et al., 2022).
@@
As discussed in Sec. 3, SMDSs can provide a solution to the enigmatic Blue Monsters uncovered by JWST. Those are extremely compact ($r_{e}\lesssim 400$pc) yet very bright objects at $z\gtrsim 10$, which show very little dust attenuation and, if interpreted as galaxies have to have converted gas to stars at incredibly high rates. Compactness, lack of dust attenuation, and high conversion rate of gas to stellar material are, in fact, a feature of the SMDS model, thus turning an astronomical puzzle into a potential line of evidence for SMDSs. Furthermore, We argue that Dark Stars can also be part of the solution to the Little Red Dots (LDRs) puzzle (e.g. Matthee et al., 2024; Kokorev et al., 2024; Akins et al., 2025). Those are extremely compact ($r_{eff}\lesssim 200$ pc) sources unlikely to be galaxies due to incredibly dense stellar packing and little to no obscuration due to dust (Casey et al., 2025). Instead many LRDs seem to point towards the existence of SMBHs surrounded by remnant stellar layers that explain the weak UV and lack of X-ray emissions (e.g. Naidu et al., 2025; de Graaff et al., 2025; Begelman and Dexter, 2025). Such objects, a.k.a. Black Hole Stars or “quasi-stars” are hypothesized in the DCBH scenario (Begelman et al., 2006) and in future work we plan to investigate if such quasi-stars they can also form via the collapse of SMDSs. In addition, a large sub-sample of JWST’s Little Red Dots (LRDs) are thought to harbor BHs that are significantly overmassive relative to their host galaxies (Durodola et al., 2025). Each of those Over-Massive Black Hole Galaxies (OBGs) can be easily explained by Dark Stars via the one of the two mechanisms we proposed in this paper for UHZ1: mergers (Sec. 5.1) or same halo (Sec. 5.2).
Moreover, Dark Stars could affect the re-ionization history of the Universe (Scott et al., 2011). As shown by Gondolo et al. (2022), the inclusion of a population of DSs can naturally lead values of CMB electron scattering optical depth as low as $\tau\sim 0.05$, in agreement with the Planck measurement. This value is impossible to reconcile with only a Pop III/II stellar population responsible for reionization, unless either very low star formation efficiencies, or strong LW feedback are invoked. In contrast, in view of their small predicted abundance, Direct Collapse Black Holes are not expected to contribute significantly to re-ionization history or to its CMB markers.
@@ Appendix A Dark Matter, a brief overview
Since Dark Stars rely on Dark Matter particles as a power engine, or, conversely, as they can be used as DM probes, we briefly review here the main evidence for the existence of Dark Matter (DM) and strategies for its eventual discovery. The fundamental nature of DM has remained largely elusive since Fritz Zwicky coined the term Dunkle Materie (i.e. Dark Matter) in 1933 (Zwicky, 1933). Its gravitational effects have since been inferred experimentally at scales ranging from galactic (e.g. Rubin and Ford, 1970), galactic cluster (e.g. Natarajan et al., 2017), and cosmological (e.g. Madhavacheril et al., 2015; Vikram et al., 2015; Hikage et al., 2019; Planck Collaboration et al., 2020). Moreover, DM plays a significant role in the early universe, leaving its imprint in afterglow radiation from the Big Bang, the so-called Cosmic Microwave Background (CMB) radiation. A combination of all those data leads to the emergence of the concordance cosmological model ($\Lambda$-CDM), where the vast majority of the energy budget of the Universe is in the form of two mysterious components: Dark Energy ($70\%$) and Dark Matter (25%). The former is responsible for the current accelerated expansion of our universe (Riess et al., 1998), while the latter assembles in large scale filamental structures, thus providing a scaffolding for the ordinary matter to assemble in galaxies and clusters. The thought that all the stars in all the galaxies in our universe form less than one percent of the total energy budget of our Universe is unsettling, perhaps even scary.
@@
It is now firmly established that the majority of the DM in the Universe comes in the form of new, yet to be discovered particles. Amongst those, Axions and Weakly Interacting Massive Particles (WIMPs) are perhaps the two leading candidates, in view of their natural theoretical motivations. There are three different strategies with which one can approach the hunt for Dark Matter particles: direct detection experiments (XENON (e.g. Aprile and et al., 2018; Aprile et al., 2020), LZ (e.g. Aalbers et al., 2023, 2024), PICO (e.g. Amole et al., 2019), etc), particle production in accelerators (e.g. Boveia and Doglioni, 2018) and lastly, indirect detection. The later method relies on potential observable signals of DM self-annihilating, or decaying, in various astrophysical objects or environments. As of yet, none of those methods lead to a discovery signal, although there are tantalizing hints, especially from indirect detection methods. For example, both the AMS-02 anomalous anti-proton signal, and the Fermi Gamma ray excess from the Galactic Center, can both be explained by the self annihilation of $\sim 60$ GeV WIMPs (Cholis et al., 2019). For both of those excesses there are also non-DM astrophysical interpretations. This is possible because of the somewhat large uncertainties, at a theoretical level, in the backgrounds for both of those signals. For reviews on Dark Matter and its theoretical or observational status the reader can consult Refs. (Freese, 2017; Arbey and Mahmoudi, 2021).
@@
We end this section with a non exhaustive list of objects that can act as indirect DM probes: neutron stars (e.g. Bramante et al., 2017; Baryakhtar et al., 2017; Bell et al., 2018, 2020; Ilie et al., 2020; Nguyen and Tait, 2023), white dwarfs (e.g. Dasgupta et al., 2019; Horowitz, 2020), Jovian planets (e.g. Leane and Smirnov, 2021; Croon and Smirnov, 2023), present day stars capturing and burning Dark Matter (e.g. Moskalenko and Wai, 2007; Scott et al., 2009; John et al., 2024), and, of particular interest for this work, the first generation of stars in the Universe (e.g. Freese et al., 2008b; Iocco, 2008; Ilie et al., 2021; Ilie and Levy, 2021).